Owusu is six times as old as kwame. In twelve years Owusu will be two times kwame. How old are they?
Answer by stanbon(75887)

Owusu is six times as old as kwame. In twelve years Owusu will be two times kwame. How old are they?
----
Equations:
O = 6*K
O+12 = 2(K+12)
----------------
Substitute for "O" and solve for "K"::
6K+12 = 2K+24
4K = 12
K = 3 yrs (Kwame's age now)
----
O = 6*K = 18 yrs (Owusu's age now)
-------------
Cheers,
Stan H.
------------